# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 3.7 Foliations attached to Hodge loci
In this section, we write the expression of a foliation in ${\sf T}:=F^{\frac{{m}}{2}}H^{m}_{\rm dR}(Y/V)$ with Hodge loci as leaf schemes. A version of these foliations developed in [Mov22, Chapter 5,6] is not suitable for our purpose and we mainly use the version in [Mov17b].
@@
(34) $x=\begin{pmatrix}0\\ \vdots\\ 0\\ x^{\frac{{m}}{2}}\\ \vdots\\ x^{{m}}\end{pmatrix}.$
Here, $x^{i}$ is a ${\sf h}^{{m}-i,i}\times 1$ matrix. We take ${\sf C}$ the constant matrix which is obtained by replacing $x_{i}$ with 00 in $x$ except for $x_{1}$, which is replaced with $1$ (this is the first coordinate of $x^{\frac{{m}}{2}}$). Let ${\sf S}$ be a Hodge block lower triangular ${\sf h}\times{\sf h}$ matrix which is obtained from the identity matrix by replacing the ${\sf h}^{\frac{{m}}{2}+1}+1$ column with $x$. It is defined in this way to have the equality
(35) ${\sf S}\cdot{\sf C}=x.$
In this way ${\sf S}^{-1}$ is obtained from ${\sf S}$ by replacing $x_{1}$ with $x_{1}^{-1}$ and $x_{i},\ i\geq 2$ with $-x_{i}x_{1}^{-1}$. Note that $\det({\sf S})=x_{1}$. Define
@@
(36) $\alpha:={\sf S}^{-1}\cdot\omega.$
Let $\nabla:H^{m}_{\rm dR}(Y/{V})\to\Omega_{V}\otimes_{{\cal O}_{V}}H^{m}_{\rm dR}(Y/{V})$ be the algebraic Gauss-Manin connection. We can write $\nabla$ in the basis $\omega$ and define the ${\sf h}\times{\sf h}$ matrix ${\sf B}$ by the equality:
@@
The entries of ${\sf B}$ are differential 1-forms in ${V}$. In a similar way we can compute the Gauss-Manin connection of ${\sf X}/{\sf T}$ in the basis $\alpha$:
(37) $\nabla\alpha={\sf A}\otimes\alpha,\ \ {\sf A}=-{\sf S}^{-1}d{\sf S}+{\sf S}^{-1}\cdot{\sf B}\cdot{\sf S}.$
This follows from the construction of the global sections $\alpha$ in (36) and the Leibniz rule. We call ${\sf B}$ (resp. ${\sf A}$) the Gauss-Manin connection matrix of the pair $(Y/{V},\omega)$ (resp. $({\sf X}/{\sf T},\alpha)$). From the integrability of the Gauss-Manin connection it follows that
(38) $d{\sf A}={\sf A}\wedge{\sf A}.$
@@
The entries of ${\sf A}{\sf C}$ induce a holomorphic foliation ${\cal F}({\sf C})$ in ${\sf T}$. The integrability follows from (38):
@@
The foliation ${\cal F}({\sf C})$ in ${\sf T}$ is given by
(39) 00 0 $=$ ${\sf B}^{\frac{{m}}{2}-1,\frac{{m}}{2}}x^{\frac{{m}}{2}}$
(40) $dx^{\frac{{m}}{2}}$ $=$ ${\sf B}^{\frac{{m}}{2},\frac{{m}}{2}}x^{\frac{{m}}{2}}+{\sf B}^{\frac{{m}}{2},\frac{{m}}{2}+1}x^{\frac{{m}}{2}+1},$
(41) $dx^{i}$ $=$ $\sum_{j=\frac{{m}}{2}}^{m}{\sf B}^{i,j}x^{j},\ \ \ \ \ i=\frac{{m}}{2}+1,\ldots,{m}.$
@@
For this we use (37) and we conclude that ${\cal F}({\sf C})$ is given by $(-{\sf S}^{-1}d{\sf S}+{\sf S}^{-1}\cdot{\sf B}\cdot{\sf S}){\sf C}$. Since ${\sf C}$ is a constant vector and ${\sf S}$ is an invertible matrix and we have (35), we conclude that ${\cal F}({\sf C})$ is given by the entries of
@@
Opening this equality and using the zero blocks of $x$ in (35) we get (39), (40) and (41). Note that by Griffiths transversality ${\sf B}^{i,j}=0$ for $j-i\geq 2$. ∎
Let $\delta_{t}\in H_{m}({\sf X}_{t},\mathbb{Q})\otimes_{\mathbb{Q}}\mathbb{C},\ t\in({\sf T},0)$ be a continuous family of cycles, that is, the Poincaré dual of $\delta_{t}$ is a flat section of the Gauss-Manin connection: $\nabla\delta_{t}=0$. Here, $({\sf T},0)$ is a small neighborhood of 00 in ${\sf T}$ in the usual topology.
@@
(42) $L_{\delta_{t}}:=\left\{\ \ \ t\in({\sf T},0)\ \ \ |\ \ \ \int_{\delta_{t}}\alpha={\sf C}.\right\},$
is a leaf scheme of ${\cal F}({\sf C})$. In other words, the ideal ${{\cal I}}_{\delta_{t}}$ generated by the entries of $\int_{\delta_{t}}\alpha-{\sf C}$ gives a leaf scheme of ${\cal F}({\sf C})$. Moreover, $\delta_{t}$ is a general Hodge cycle if and only if 00 is a smooth point of ${\cal F}(\Omega)$, and hence, $L_{\delta_{t}}$ is general.
@@
(43) $df=\int_{\delta_{t}}\nabla\alpha={\sf A}\cdot\int_{\delta_{t}}\alpha={\sf A}\cdot{\sf C}+{\sf A}\cdot f.$
This implies that $L_{\delta_{t}}$ is a leaf scheme of ${\cal F}({\sf C})$. For the second part, note that ${\sf B}^{\frac{{m}}{2}-1,\frac{{m}}{2}}x^{\frac{{m}}{2}}$ is just the matrix format of $\alpha$ in (29), see [Mov17b, Proposition 5]. The singular locus of ${\cal F}({\sf C})$ corresponds to those $t\in{\sf T}$ such that the kernel of the entries of ${\sf B}^{\frac{{m}}{2}-1,\frac{{m}}{2}}x^{\frac{{m}}{2}}$ is not minimal, and the statement follows. ∎
One might conjecture that all the leaf schemes of ${\cal F}({\sf C})$ are of the form $L_{\delta_{t}}$. As we do not need this kind of statements, we leave it to the reader.
@@
The Hodge locus with constant periods ${\sf C}$ is defined to be $L_{\delta_{t}}$ in (42) with $\delta_{t}\in H^{m}({\sf X}_{t},\mathbb{Q})$. Its ideal is given by
${{\cal I}}_{\delta_{t}}:=\left\langle\int_{\delta_{t}}\alpha-{\sf C}\right\rangle=\left\langle\int_{\delta_{t}}\omega-x\right\rangle\subset{\cal O}_{{\sf T}^{\rm hol},0}.$
@@
For the example mentioned in the Introduction, we take $E_{z}:y^{2}=x(x-1)(x-t)$ the Legendre family of elliptic curves, and $Y_{t_{1},t_{2}}:=E_{t_{1}}\times E_{t_{2}}$. The de Rham cohomology $H^{2}_{\rm dR}(Y_{t_{1},t_{2}})$ modulo the cohomology class of fibers of projections in each factor, is four dimensional with Hodge numbers $1,2,1$. The loci of isogenies of degree $N$ between $E_{t_{1}}$ and $E_{t_{2}}$ is an algebraic curve and we can construct the corresponding leaf scheme and foliation in $(t_{1},t_{2},x_{1},x_{2},x_{3})\in\mathbb{C}^{5}$. For a more conceptual treatment of this example see [Mov22, Chapter 10].
@@ 3.8 The definition ring of Hodge loci
Let $Y\to{V}$ be a family of smooth projective varieties over $\mathbb{C}$ and take a smooth model of this over ${\mathfrak{R}}\subset\mathbb{C}$. Let also 00 be an ${\mathfrak{R}}$-valued point of ${V}$. We take a topological cycle $\delta_{0}\in H_{m}(Y_{0},\mathbb{Z})$ and enlarge ${\mathfrak{R}}$ to
@@
that is, ${\mathfrak{R}}$ is the ring of polynomials in the periods $(2\pi i)^{-\frac{m}{2}}\int_{\delta_{0}}\omega,\omega\in H^{m}_{\rm dR}(Y_{0}/{\mathfrak{R}})$. Note that these numbers are conjecturally in $\bar{\mathfrak{k}}$, where ${\mathfrak{k}}$ is the quotient field of ${\mathfrak{R}}$. This follows from the Hodge conjecture or if we assume that $\delta_{0}$ is an absolute Hodge cycle. We will not need this for our investigation, and hence, ${\mathfrak{R}}(\delta_{0})$ might contain new transcendental numbers. Consider the monodromy $\delta_{t}\in H_{m}(Y_{t},\mathbb{Z})$ of $\delta_{0}$ to nearby fibers $t\in({V},0)$.
@@
Let 00 be an ${\mathfrak{R}}$-valued point of ${V}$ and $\omega\in H^{m}_{\rm dR}(Y/{V})$. Then the Taylor series of $\int_{\delta_{t}}\omega$ at 00 has coefficients in ${\mathfrak{R}}(\delta_{0})_{\mathbb{Q}}:={\mathfrak{R}}(\delta_{0})\otimes_{\mathbb{Z}}\mathbb{Q}$, that is,
$\int_{\delta_{t}}\omega\in{\cal O}_{{V}^{\rm hol}_{{\mathfrak{R}}(\delta_{0})_{\mathbb{Q}}},0}.$
@@
This follows from 3.3 and 3.2. The main ingredient is that the Gauss-Manin connection matrix of of $Y/{V}$ is defined over ${\mathfrak{R}}$. ∎
It might be too naive to believe that if $\delta_{0}$ is a Hodge cycle then it has coefficients in ${\mathfrak{R}}(\delta_{0})$ itself, that is, $\int_{\delta_{t}}\omega\in{\cal O}_{{V}^{\rm hol}_{{\mathfrak{R}}(\delta_{0})},0}$. In Section 3.10 we experimentally observe that this is false.
@@
The Cattani-Deligne-Kaplan theorem, see [CDK95, Theorem 1.1], implies that there is an algebraic subvariety $L$ of ${\sf T}_{\mathbb{C}}$ such that $L_{\delta_{t}}$ is just an analytic germ/small open subset of $L$. By monodromy of $\delta_{t}$ along any path in $L$, we can see that $L$ is covered by such open sets. Therefore, $L$ has the structure of an analytic scheme given by analytic ideals ${{\cal I}}_{\delta_{t}}$. We have to consider a compactification $\bar{\sf T}_{\mathbb{C}}$ and see that the analytic scheme structure of $L$ extends to $\bar{L}$. This is not explicitly mentioned in [CDK95], but their proof implies this. Now by GAGA for analytic subschemes of projective varieties we conclude that ${{\cal I}}_{\delta_{t}}$ is algebraic. 3.1 finishes the proof. ∎
@@
Let ${\mathfrak{k}}$ be the quotient field of ${\mathfrak{R}}$. If $\delta_{0}$ is an absolute Hodge cycle then all its periods $\frac{1}{(2\pi i)^{\frac{m}{2}}}\int_{\delta_{0}}\omega,\omega\in H^{m}_{\rm dR}(Y_{0})$ are in the algebraic closure $\bar{\mathfrak{k}}$ of ${\mathfrak{k}}$. Since ${\cal F}({\sf C})$ is defined over ${\mathfrak{k}}$, by taking the action of the Galois group ${\rm Gal}(\bar{\mathfrak{k}}/{\mathfrak{k}})$ on the coefficients of $L$ and the base point $(0,x_{0})$, we get finitely many leaf schemes of ${\cal F}(\Omega)$ which come from Hodge cycles too. If $\delta_{0}$ is not absolute then some of the period of $\delta_{0}$ are transcendental numbers, that it they do not belong to $\bar{\mathfrak{k}}$. In this case by some standard arguments, see for instance [Mov22, Theorem 5.18], we can transform such transcendental numbers into variables and get continuous families of algebraic leaf schemes for ${\cal F}({\sf C})$. In general it is open whether such leaf schemes come from Hodge cycles or not. For some related results see [Mov22, Section 7.5].

### v2

@@ 3.7 Foliations attached to Hodge loci
In this section, we write the expression of a foliation in ${\sf T}:=F^{\frac{{m}}{2}}H^{m}_{\rm dR}(Y/V)$ with Hodge loci as leaf schemes. A version of these foliations developed in [ho2020, Chapter 5,6] is not suitable for our purpose and we mainly use the version in [GMCD-NL].
@@
(3.8) $x=\begin{pmatrix}0\\ \vdots\\ 0\\ x^{\frac{{m}}{2}}\\ \vdots\\ x^{{m}}\end{pmatrix}$.
Here, $x^{i}$ is a ${\sf h}^{{m}-i,i}\times 1$ matrix. We take $\sf C$} the constant matrix which is obtained by replacing $x_{i}$ with 0 in $x$ except for $x_{1}$, which is replaced with $1$ (this is the first coordinate of $x^{\frac{{m}}{2}}$). Let $\sf S$} be a Hodge block lower triangular ${\sf h}\times{\sf h}$ matrix which is obtained from the identity matrix by replacing the ${\sf h}^{\frac{{m}}{2}+1}+1$ column with $x$. It is defined in this way to have the equality
(3.9) ${\sf S}\cdot{\sf C}=x.$
In this way ${\sf S}^{-1}$ is obtained from $\sf S$} by replacing $x_{1}$ with $x_{1}^{-1}$ and $x_{i},\ i\geq 2$ with $x_{i}x_{1}^{-1}$. Note that $\det({\sf S})=x_{1}$. Define
@@
(3.10) $\alpha:={\sf S}^{-1}\cdot\omega.$
Let $\nabla:H^{m}_{\rm dR}(Y/{V})\to\Omega_{V}\otimes_{{\mathcal{O}}_{V}}H^{m}_{\rm dR}(Y/{V})$ be the algebraic Gauss-Manin connection. We can write $\nabla$ in the basis $\omega$ and define the ${\sf h}\times{\sf h}$ matrix $\sf B$} by the equality:
@@
The entries of $\sf B$} are differential 1-forms in V{V}. In a similar way we can compute the Gauss-Manin connection of ${\sf X}/{\sf T}$ in the basis $\alpha$:
(3.11) $\nabla\alpha={\sf A}\otimes\alpha,\ \ {\sf A}=-{\sf S}^{-1}d{\sf S}+{\sf S}^{-1}\cdot{\sf B}\cdot{\sf S}.$
This follows from the construction of the global sections $\alpha$ in (3.10) and the Leibniz rule. We call $\sf B$} (resp. $\sf A$}) the Gauss-Manin connection matrix of the pair $(Y/{V},\omega)$ (resp. $({\sf X}/{\sf T},\alpha)$). From the integrability of the Gauss-Manin connection it follows that
(3.12) $d{\sf A}={\sf A}\wedge{\sf A}.$
@@
The entries of ${\sf A}{\sf C}$ induce a holomorphic foliation ${\mathcal{F}}({\sf C})$ in $\sf T$}. The integrability follows from (3.12):
@@
The foliation ${\mathcal{F}}({\sf C})$ in $\sf T$} is given by
(3.13) $0$ $=$ ${\sf B}^{\frac{{m}}{2}-1,\frac{{m}}{2}}x^{\frac{{m}}{2}}$
(3.14) $dx^{\frac{{m}}{2}}$ $=$ ${\sf B}^{\frac{{m}}{2},\frac{{m}}{2}}x^{\frac{{m}}{2}}+{\sf B}^{\frac{{m}}{2},\frac{{m}}{2}+1}x^{\frac{{m}}{2}+1},$
(3.15) $dx^{i}$ $=$ $\sum_{j=\frac{{m}}{2}}^{m}{\sf B}^{i,j}x^{j},\ \ \ \ \ i=\frac{{m}}{2}+1,\ldots,{m}.$
@@
For this we use (3.11) and we conclude that ${\mathcal{F}}({\sf C})$ is given by $(-{\sf S}^{-1}d{\sf S}+{\sf S}^{-1}\cdot{\sf B}\cdot{\sf S}){\sf C}$. Since $\sf C$} is a constant vector and $\sf S$} is an invertible matrix and we have (3.9), we conclude that ${\mathcal{F}}({\sf C})$ is given by the entries of
@@
Opening this equality and using the zero blocks of $x$ in (3.9) we get (3.13), (3.14) and (3.15). Note that by Griffiths transversality ${\sf B}^{i,j}=0$ for $j-i\geq 2$. ∎
Let $\delta_{t}\in H_{m}({\sf X}_{t},\mathbb{Q})\otimes_{\mathbb{Q}}\mathbb{C},\ t\in({\sf T},0)$ be a continuous family of cycles, that is, the Poincaré dual of $\delta_{t}$ is a flat section of the Gauss-Manin connection: $\nabla\delta_{t}=0$. Here, $({\sf T},0)$ is a small neighborhood of 0 in $\sf T$} in the usual topology.
@@
(3.16) $L_{\delta_{t}}:=\left\{\ \ \ t\in({\sf T},0)\ \ \ |\ \ \ \int_{\delta_{t}}\alpha={\sf C}.\right\},$
is a leaf scheme of ${\mathcal{F}}({\sf C})$. In other words, the ideal ${{\mathcal{I}}}_{\delta_{t}}$ generated by the entries of $\int_{\delta_{t}}\alpha-{\sf C}$ gives a leaf scheme of ${\mathcal{F}}({\sf C})$. Moreover, $\delta_{t}$ is a general Hodge cycle if and only if 0 is a smooth point of ${\mathcal{F}}(\Omega)$, and hence, $L_{\delta_{t}}$ is general.
@@
(3.17) $df=\int_{\delta_{t}}\nabla\alpha={\sf A}\cdot\int_{\delta_{t}}\alpha={\sf A}\cdot{\sf C}+{\sf A}\cdot f.$
This implies that $L_{\delta_{t}}$ is a leaf scheme of ${\mathcal{F}}({\sf C})$. For the second part, note that ${\sf B}^{\frac{{m}}{2}-1,\frac{{m}}{2}}x^{\frac{{m}}{2}}$ is just the matrix format of $\alpha$ in (3.3), see [GMCD-NL, Proposition 5]. The singular locus of ${\mathcal{F}}({\sf C})$ corresponds to those $t\in{\sf T}$ such that the kernel of the entries of ${\sf B}^{\frac{{m}}{2}-1,\frac{{m}}{2}}x^{\frac{{m}}{2}}$ is not minimal, and the statement follows. ∎
One might conjecture that all the leaf schemes of ${\mathcal{F}}({\sf C})$ are of the form $L_{\delta_{t}}$. As we do not need this kind of statements, we leave it to the reader.
@@
The Hodge locus with constant periods $\sf C$} is defined to be $L_{\delta_{t}}$ in (3.16) with $\delta_{t}\in H^{m}({\sf X}_{t},\mathbb{Q})$. Its ideal is given by
${{\mathcal{I}}}_{\delta_{t}}:=\left\langle\int_{\delta_{t}}\alpha-{\sf C}\right\rangle=\left\langle\int_{\delta_{t}}\omega-x\right\rangle\subset{\mathcal{O}}_{{\sf T}^{\rm hol},0}.$
@@
For the example mentioned in the Introduction, we take $E_{z}:y^{2}=x(x-1)(x-t)$ the Legendre family of elliptic curves, and $Y_{t_{1},t_{2}}:=E_{t_{1}}\times E_{t_{2}}$. The de Rham cohomology $H^{2}_{\rm dR}(Y_{t_{1},t_{2}})$ modulo the cohomology class of fibers of projections in each factor, is four dimensional with Hodge numbers $1,2,1$. The loci of isogenies of degree $N$ between $E_{t_{1}}$ and $E_{t_{2}}$ is an algebraic curve and we can construct the corresponding leaf scheme and foliation in $(t_{1},t_{2},x_{1},x_{2},x_{3})\in\mathbb{C}^{5}$. For a more conceptual treatment of this example see [ho2020, Chapter 10].
@@ 3.8 The definition ring of Hodge loci
Let $Y\to{V}$ be a family of smooth projective varieties over $\mathbb{C}$ and take a smooth model of this over $\mathfrak{R}$$\subset\mathbb{C}$. Let also 0 be an $\mathfrak{R}$}-valued point of V{V}. We take a topological cycle $\delta_{0}\in H_{m}(Y_{0},\mathbb{Z})$ and enlarge $\mathfrak{R}$} to
@@
that is, ${\mathfrak{R}}(\delta_{0})$ is the ring of polynomials in the periods $(2\pi i)^{-\frac{m}{2}}\int_{\delta_{0}}\omega,\omega\in H^{m}_{\rm dR}(Y_{0}/{\mathfrak{R}})$. Note that these numbers are conjecturally in $\bar{\mathfrak{k}}$, where $\mathfrak{k}$} is the quotient field of $\mathfrak{R}$}. This follows from the Hodge conjecture or if we assume that $\delta_{0}$ is an absolute Hodge cycle. We will not need this for our investigation, and hence, ${\mathfrak{R}}(\delta_{0})$ might contain new transcendental numbers. Consider the monodromy $\delta_{t}\in H_{m}(Y_{t},\mathbb{Z})$ of $\delta_{0}$ to nearby fibers $t\in({V},0)$.
@@
Let 0 be an $\mathfrak{R}$}-valued point of V{V} and $\omega\in H^{m}_{\rm dR}(Y/{V})$. Then the Taylor series of $\int_{\delta_{t}}\omega$ at 0 has coefficients in ${\mathfrak{R}}(\delta_{0})_{\mathbb{Q}}:={\mathfrak{R}}(\delta_{0})\otimes_{\mathbb{Z}}\mathbb{Q}$, that is,
$\int_{\delta_{t}}\omega\in{\mathcal{O}}_{{V}^{\rm hol}_{{\mathfrak{R}}(\delta_{0})_{\mathbb{Q}}},0}.$
@@
This follows from 3.3 and 3.2. The main ingredient is that the Gauss-Manin connection matrix of of Y/VY/{V} is defined over $\mathfrak{R}$}. ∎
It might be too naive to believe that if $\delta_{0}$ is a Hodge cycle then it has coefficients in ${\mathfrak{R}}(\delta_{0})$ itself, that is, $\int_{\delta_{t}}\omega\in{\mathcal{O}}_{{V}^{\rm hol}_{{\mathfrak{R}}(\delta_{0})},0}$. In Section 3.10 we experimentally observe that this is false.
@@
The Cattani-Deligne-Kaplan theorem, see [cadeka, Theorem 1.1], implies that there is an algebraic subvariety $L$ of $\sf T$$_{\mathbb{C}}$ such that $L_{\delta_{t}}$ is just an analytic germ/small open subset of $L$. By monodromy of $\delta_{t}$ along any path in $L$, we can see that $L$ is covered by such open sets. Therefore, $L$ has the structure of an analytic scheme given by analytic ideals ${{\mathcal{I}}}_{\delta_{t}}$. We have to consider a compactification $\bar{\sf T}_{\mathbb{C}}$ and see that the analytic scheme structure of $L$ extends to $\bar{L}$. This is not explicitly mentioned in [cadeka], but their proof implies this. Now by GAGA for analytic subschemes of projective varieties we conclude that ${{\mathcal{I}}}_{\delta_{t}}$ is algebraic. 3.1 finishes the proof. ∎
@@
Let $\mathfrak{k}$} be the quotient field of $\mathfrak{R}$}. If $\delta_{0}$ is an absolute Hodge cycle then all its periods $\frac{1}{(2\pi i)^{\frac{m}{2}}}\int_{\delta_{0}}\omega,\omega\in H^{m}_{\rm dR}(Y_{0})$ are in the algebraic closure $\bar{\mathfrak{k}}$ of $\mathfrak{k}$}. Since ${\mathcal{F}}({\sf C})$ is defined over $\mathfrak{k}$}, by taking the action of the Galois group ${\rm Gal}(\bar{\mathfrak{k}}/{\mathfrak{k}})$ on the coefficients of $L$ and the base point $(0,x_{0})$, we get finitely many leaf schemes of ${\mathcal{F}}(\Omega)$ which come from Hodge cycles too. If $\delta_{0}$ is not absolute then some of the period of $\delta_{0}$ are transcendental numbers, that it they do not belong to $\bar{\mathfrak{k}}$. In this case by some standard arguments, see for instance [ho2020, Theorem 5.18], we can transform such transcendental numbers into variables and get continuous families of algebraic leaf schemes for ${\mathcal{F}}({\sf C})$. In general it is open whether such leaf schemes come from Hodge cycles or not. For some related results see [ho2020, Section 7.5].